OpenStudy (anonymous):

Johnny makes wooden benches to sell. His advertising costs are $150 per month and each bench costs him $100 to make. If he sells each bench for $175, how many benches must he sell to break even each month?

OpenStudy (anonymous):

he needs to at least sell 2 benches to make a profit

OpenStudy (kagıtucak):

lets say he needs to sell x branches then he will consume for a month 150+x.100$ and he will sell it for 175$then he will get 175x $ for a month. If 150+100x=175x he will break 150=75x x=2

OpenStudy (anonymous):

2 or 1 depending on how many months

OpenStudy (kagıtucak):

question says each month so 2

OpenStudy (anonymous):

150 + 100x = 175x 150 = 175x - 100x 150 = 75x 150/75 = x 2 = x I agree...he has to sell 2 benches per month to break even